# Democratizing Agentic AI with Fast Test-Time Scaling on the Edge

### v1

@@ Abstract.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
(b) Reduced latency by FlashTTS
Figure 1. (a) Memory cost across models. (b) FlashTTS enables low-latency edge deployment of reasoning models. Cloud accuracy: GPT-o1-preview. Edge accuracy: Qwen2.5-Math-1.5B. Cloud latency from the first-answer latency of GPT-o3-pro and GPT-5 (thinking models) (art, 2025; Yang et al., 2025b).
[…]
Recent advances in reasoning LLMs have unlocked significant progress in solving complex tasks such as multi-hop question answering, tool use, and long-horizon planning (Google DeepMind, 2025; Team et al., 2025; Yang et al., 2025a). These capabilities are foundational for agentic AI systems, where AI agents can plan, act, and interact autonomously. As such systems move closer to real-world deployment, there is a growing demand to deploy strong reasoning LLMs at the edge (e.g., on AI PCs), where agentic systems can preserve data privacy, enable personalization, operate offline or with limited connectivity, and interact with local environments using high-level intelligence. However, edge hardware imposes severe memory limitations (e.g., a single consumer GPU with 8–24 GB VRAM), restricting deployment to edge LLMs ($\leq$ 7B) that cannot match the reasoning performance of large cloud models, limiting their effectiveness in complex tasks. As shown in Fig. 1(a), the memory capacity of consumer-grade GPUs restricts deployment to models like Qwen2.5-Math-1.5B, resulting in a significant gap in reasoning ability compared to large-scale cloud LLMs.
Deploying strong reasoning LLMs on the edge is essential for realizing democratized agentic AI, where intelligent agents are decentralized and run directly on client-side devices for better privacy and local integration. To achieve this, Test-Time Scaling (TTS) (Snell et al., 2024; Beeching et al., [n. d.]) has recently emerged as a promising candidate to bridge the reasoning gap between small, edge-deployable LLMs and large cloud-based models. Instead of relying on scaling model parameters during training, TTS allocates additional compute during inference to improve generation and reasoning quality. Despite its great potential, deploying TTS naively on existing serving systems incurs significant latency overhead, making it impractical for real-time applications. As shown in  Fig. 1(b), using a baseline vLLM implementation to match the accuracy of a large cloud model results in 200 seconds of latency, nearly doubling the latency of large models on cloud infrastructure. To realize the vision of democratized agentic AI, there is an urgent need for an efficient, edge-ready serving infrastructure that makes TTS both performant and practical.
@@
To overcome these obstacles, we present FlashTTS, a serving system that integrates three synergistic optimizations to make TTS practical on edge devices. To address Challenge-1, we introduce Speculative Beam Extension that generates speculatively to hide the latency of irregular workloads. To tackle Challenge-2 and Challenge-3, FlashTTS combines two memory-aware optimizations: Dynamic Prefix-Aware Scheduling reorders execution to maximize KV cache reuse from dynamic prefix sharing, and Asymmetric Multi-Model Memory Allocation intelligently partitions memory between the generator and verifier to improve throughput. Together, we push the boundaries of edge deployment of TTS (Fig. 1(b)), making fast and high-quality reasoning feasible on memory-constrained edge devices.
@@
  • We design and implement FlashTTS, a high-performance serving system for TTS that incorporates three novel and synergistic optimizations: Speculative Beam Extension, Dynamic Prefix-Aware Scheduling, and Asymmetric Multi-Model Memory Allocation.
@@ 2.1. LLM Reasoning
Reasoning is a critical capability for Large Language Models (LLMs), enabling multi-step problem solving and complex decision-making. This reasoning capability is initially established through reinforcement learning methods such as Guided Reinforcement Policy Optimization (GRPO), as exemplified by DeepSeek-R1 (Team et al., 2025). Such RL training fosters emergent abilities like long Chain-of-Thought (CoT) reasoning, which in turn expands the applicability of LLMs to domains including mathematical problem solving (Shao et al., 2024), scientific discovery (Google DeepMind, 2025; Mirza et al., 2025; Yang et al., 2024), coding assistant (Guo et al., 2024), and multi-hop question answering (Yang et al., 2025a; Team et al., 2025).
@@ 2.2. Test-Time Scaling (TTS) Methods
While long Chain-of-Thought (CoT) reasoning enhances the capabilities of LLMs, smaller models still lag significantly behind their larger counterparts (Liu et al., 2025). To bridge this gap, Test-Time Scaling (TTS) increases the computational budget for inference by exploring multiple reasoning paths in parallel (Snell et al., 2024; Beeching et al., [n. d.]; Liu et al., 2025). Early TTS methods mainly relied on Best-of-N (BoN) sampling, where an Outcome Reward Model (ORM) selects the best solution from a set of fully generated candidates (Wang et al., 2022; Cobbe et al., 2021). However, BoN offers limited guidance during generation and yields less diversity in reasoning path structures ( Fig. 2). The introduction of Process Reward Models (PRMs), which evaluate intermediate reasoning steps, has enabled advanced verifier-guided search algorithms such as Beam Search and DVTS (Lightman et al., 2023; Uesato et al., 2022; Snell et al., 2024). These methods follow a generation–verification paradigm: a PRM periodically scores partial solutions, expanding high-scoring trajectories and pruning weak ones, thereby concentrating computation on promising paths (Hooper et al., 2025; Wu et al., 2024b; Chen et al., 2025). As a result, the LLM produces a diverse reasoning tree rather than a single chain.
PRMs are primarily categorized as either discriminative or generative (Malik et al., 2025). A discriminative PRM functions as a sequence classifier; in a single forward pass, it takes a full reasoning path as input and outputs a score for each intermediate step (He et al., 2024; Wang et al., 2023; Lightman et al., 2023). In contrast, a generative PRM is an auto-regressive model that must first generate its own textual critique before providing a final score, a significantly more expensive process (Zhao et al., 2025). Due to their superior balance of model parameters, reasoning quality, and hardware efficiency, discriminative PRMs are the preferred choice for state-of-the-art, verifier-guided TTS systems, particularly for memory-constrained edge deployment (Malik et al., 2025). Hence, our system focuses on discriminative PRMs. In contrast, as noted by (Snell et al., 2024; Liu et al., 2025), multi-step lookahead approaches (Chen and Li, 2024; Wang et al., 2025), such as Monte Carlo Tree Search (MCTS) (Feng et al., 2023), introduce significant sampling and latency overhead with inferior accuracy, hence we do not consider them in this work.
@@ 2.3. LLM Serving
Serving frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024b) have been developed to optimize throughput and latency in streaming query scenarios. These systems incorporate key optimizations, including KV cache management to avoid recomputing attention states, paged attention to reduce GPU memory fragmentation, and preemptive scheduling to handle memory constraints by swapping requests. For TTS serving, goodput will be a more useful metric rather than throughput, as not all generated tokens will be selected for the final output. Despite its importance, no serving system to date natively supports the structured, multi-path search required for TTS in reasoning tasks.
@@ 3.1. Patterns Analysis in TTS Methods
Recent advancements in LLM reasoning have led to a variety of TTS methods, evolving from simple parallel sampling to more sophisticated search methods (Zhang et al., 2025; Chen et al., 2025; Snell et al., 2024; Hooper et al., 2025; Beeching et al., [n. d.]). As illustrated in Fig. 2, this evolution marks a structural shift in the generation process: from exploring parallel but independent chains (e.g., CoT and Best-of-N) to constructing complex reasoning trees that allow for intermediate guidance and pruning (e.g., Beam Search and its variants).
@@
This two-step process repeats until all reasoning paths reach a terminal state. Various search algorithms shown in Fig. 2 can be understood as specific implementations of this general pattern, differing in the heuristics they apply during the Generation or Verification stage. For instance, during Verification, standard Beam Search selects the top-K candidates globally with a static branching factor. In contrast, Diverse Selection (Snell et al., 2024; Beeching et al., [n. d.]) modifies this to improve diversity by choosing the top candidate from distinct subtrees, while Dynamic Branching (Hooper et al., 2025; Wu et al., 2024b) makes the branching factor itself adaptive to verifier scores and system state. Other methods, like Varying Granularity (Chen et al., 2025), instead modify the Generation stage by altering the length of the thinking steps produced between verifications.
To understand the accuracy-latency trade-offs across different TTS methods, we conduct evaluations on the MATH-500 dataset. As illustrated in Fig. 3 (left), while advanced search methods often achieve higher algorithm accuracy, their overall latency remains a critical bottleneck. To address this system-level performance gap, we analyze the challenges shared by the abstracted TTS pattern.
@@ 3.2.3. Constrained Memory for Multi-Model Execution
4. FlashTTS: Method and Optimization
@@ 4.1.2. Two-Phase Scheduling with Preemption
The speculative phase is fully preemptible: if a new request arrives or a running request is preempted due to memory constraints, all speculative generation is immediately stopped, and the system reverts to Phase 1 to serve the new request. This two-phase design ensures minimal overhead and preserves low-latency responsiveness.
@@ 4.3.1. Roofline-Guided KV Allocation
$T_{\mathrm{roof}}=\max\!\left(\frac{\mathrm{FLOPs}}{P\cdot 10^{12}},\;\frac{\mathrm{Bytes}}{\mathrm{BW}\cdot 10^{9}}\right),$
where $P$ is the device’s peak compute (TFLOP/s) and $\mathrm{BW}$ its peak memory bandwidth (GB/s), per hardware specification.
@@ 4.3.2. Extended Search Space with Offloading
This incurs a transfer overhead $T_{\text{overhead}}^{\text{offload}}$. The system then selects the lower-latency strategy: i) the optimal execution time $T_{\mathrm{tot}}$ from allocation search under the original constraint, or ii) the offloading time $T_{\mathrm{tot}}^{\text{offload}}+T_{\text{overhead}}^{\text{offload}}$, where $T_{\mathrm{tot}}^{\text{offload}}$ is computed from the maximum batch sizes allowed by the relaxed constraints. This dual-strategy policy lets FlashTTS always pick the better option.
@@ 5. Implementation
FlashTTS is implemented in ~6,500 lines of Python on top of vLLM (v0.9.2), operating the generator and verifier in separate worker processes via Python’s multiprocessing library. We extend the core LLMEngine of vLLM to implement our two-phase, preemptive scheduling policy, which dynamically switches between continuous batching and Speculative Beam Extension based on the request queue status. For the generator, we extend the default scheduler with our Dynamic Prefix-Aware Scheduling that implements a greedy heuristic to group beams from the same parent, maximizing KV cache reuse. Our Asymmetric Multi-Model Memory Allocation policy is managed by a lightweight searcher that is invoked dynamically to determine the partition of the KV cache between workers. The system exposes a configurable interface for various TTS strategies and hyperparameters.
Figure 11. Precise Goodput improvement of FlashTTS over the vLLM baseline across different search algorithm variants. Experiments use the 1.5B+1.5B configuration on AIME. In dynamic branching, each beam branches proportionally to its verifier score; in varying granularity, the maximum step length is 64 tokens for the first 3 steps and 2048 thereafter.
Figure 12. FlashTTS Goodput Improvement. The $x$-axis represents the number of beams (n).
@@ 6.1. Experimental Setup
Models. To assess FlashTTS under diverse workloads, we evaluate three generator-verifier configurations designed to stress different system aspects, following common practice in prior work (Snell et al., 2024; Chen et al., 2025; Beeching et al., [n. d.]; Liu et al., 2025). We test a verifier-heavy setting (1.5B+7B: Qwen2.5-Math-1.5B generator with a Math-Shepherd-Mistral-7B verifier) and a generator-heavy setting (7B+1.5B: Qwen2.5-Math-7B generator with a Skywork-o1-Open-PRM-1.5B verifier), both allocate 90% of GPU memory to test throughput limits. To simulate a highly resource-limited environment, we also test a memory-constrained setting (1.5B+1.5B: a 1.5B generator and verifier) (Yang et al., 2024; He et al., 2024; Wang et al., 2023), restricting it to 40% of GPU memory.
@@
Baseline Implementation. Our baseline system is built on top of the widely-used vLLM framework (version 0.9.2). We implement a standard verifier-guided test-time search, running the generator and verifier as separate vLLM instances, with remaining details following Hugging Face’s official search-and-learn implementation (Beeching et al., [n. d.]). This baseline represents a naive but robust implementation of TTS, against which we compare the performance gains achieved by the optimizations in FlashTTS.
@@
  • Completion Time: We measure the average end-to-end time taken per completion for a probelm.

### v2

# FastTTS: Accelerating Test-Time Scaling for Edge LLM Reasoning

@@ Abstract.
machine learning system; LLM reasoning; test-time scaling (TTS); resource-constrained inference
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790161isbn: 979-8-4007-2359-9/2026/03ccs: Computer systems organization Real-time systemsccs: Computing methodologies Multi-agent planning
@@ 1. Introduction
Recent advances in reasoning LLMs have unlocked significant progress in solving complex tasks such as multi-hop question answering, tool use, and long-horizon planning (Google DeepMind, 2025; Team and others, 2025; Yang et al., 2025a). These capabilities are foundational for agentic AI systems, where AI agents can plan, act, and interact autonomously. As such systems move closer to real-world deployment, there is a growing demand to deploy strong reasoning LLMs at the edge (e.g., on AI PCs), where agentic systems can preserve data privacy, enable personalization, operate offline or with limited connectivity, and interact with local environments using high-level intelligence. However, edge hardware imposes severe memory limitations (e.g., a single consumer GPU with 8–24 GB VRAM), restricting deployment to edge LLMs ($\leq$ 7B) that cannot match the reasoning performance of large cloud models, limiting their effectiveness in complex tasks. As shown in Fig. 1(a), the memory capacity of consumer-grade GPUs restricts deployment to models like Qwen2.5-Math-1.5B, resulting in a significant gap in reasoning ability compared to large-scale cloud LLMs.
@@
(b) Reduced latency by FastTTS
Figure 1. (a) Memory cost across models. (b) FastTTS enables low-latency edge deployment of reasoning models. Cloud accuracy: GPT-o1-preview. Edge accuracy: Qwen2.5-Math-1.5B. Cloud latency from the first-answer latency of GPT-o3-pro and GPT-5 (thinking models) (3; L. Yang, Z. Yu, B. Cui, and M. Wang (2025b)).
Deploying strong reasoning LLMs on the edge is essential for realizing democratized agentic AI, where intelligent agents are decentralized and run directly on client-side devices for better privacy and local integration. Such guarantees are particularly critical in sensitive domains such as healthcare, autonomous driving, and defense, where on-device reasoning is essential for privacy, energy efficiency, or responsiveness (Nissen et al., 2025; Saad-Falcon et al., 2025). To achieve this, Test-Time Scaling (TTS) (Snell et al., 2024; Beeching et al., ) has recently emerged as a promising candidate to bridge the reasoning gap between small, edge-deployable LLMs and large cloud-based models. Instead of relying on scaling model parameters during training, TTS allocates additional compute during inference to improve generation and reasoning quality. Despite its great potential, deploying TTS naively on existing serving systems incurs significant latency overhead, making it impractical for real-time applications. As shown in  Fig. 1(b), using a baseline vLLM implementation to match the accuracy of a large cloud model results in 200 seconds of latency, nearly doubling the latency of large models on cloud infrastructure. To realize the vision of democratized agentic AI, there is an urgent need for an efficient, edge-ready serving infrastructure that makes TTS both performant and practical.
@@
To overcome these obstacles, we present FastTTS, a serving system that integrates three synergistic optimizations to make TTS practical on edge devices. To address Challenge-1, we introduce Speculative Beam Extension that generates speculatively to hide the latency of irregular workloads. To tackle Challenge-2 and Challenge-3, FastTTS combines two memory-aware optimizations: Dynamic Prefix-Aware Scheduling reorders execution to maximize KV cache reuse from dynamic prefix sharing, and Asymmetric Multi-Model Memory Allocation intelligently partitions memory between the generator and verifier to improve throughput. Together, we push the boundaries of edge deployment of TTS (Fig. 1(b)), making fast and high-quality reasoning feasible on memory-constrained edge devices.
@@
  • We design and implement FastTTS, a high-performance serving system for TTS that incorporates three novel and synergistic optimizations: Speculative Beam Extension, Dynamic Prefix-Aware Scheduling, and Asymmetric Multi-Model Memory Allocation.
@@ 2.1. LLM Reasoning
Reasoning is a critical capability for Large Language Models (LLMs), enabling multi-step problem solving and complex decision-making. This reasoning capability is initially established through reinforcement learning methods such as Guided Reinforcement Policy Optimization (GRPO), as exemplified by DeepSeek-R1 (Team and others, 2025). Such RL training fosters emergent abilities like long Chain-of-Thought (CoT) reasoning, which in turn expands the applicability of LLMs to domains including mathematical problem solving (Shao et al., 2024), scientific discovery (Google DeepMind, 2025; Mirza et al., 2025; Yang et al., 2024), coding assistant (Guo et al., 2024; Gai et al., 2025), and multi-hop question answering (Yang et al., 2025a; Team and others, 2025).
@@ 2.2. Test-Time Scaling (TTS) Methods
While long Chain-of-Thought (CoT) reasoning enhances the capabilities of LLMs, smaller models still lag significantly behind their larger counterparts (Liu et al., 2025). To bridge this gap, TTS increases the computational budget for inference by exploring multiple reasoning paths in parallel (Snell et al., 2024; Beeching et al., ; Liu et al., 2025). Early TTS methods mainly relied on Best-of-N (BoN) sampling, where an Outcome Reward Model (ORM) selects the best solution from a set of fully generated candidates (Wang et al., 2022; Cobbe et al., 2021). However, BoN offers limited guidance during generation and yields less diversity in reasoning path structures ( Fig. 2). The introduction of Process Reward Models (PRMs), which evaluate intermediate reasoning steps, has enabled advanced verifier-guided search algorithms such as Beam Search and Monte Carlo Tree Search (MCTS) (Lightman et al., 2023; Uesato et al., 2022; Snell et al., 2024). These methods follow a generation–verification paradigm: a PRM periodically scores partial solutions, expanding high-scoring trajectories and pruning weak ones, thereby concentrating computation on promising paths (Hooper et al., 2025; Wu et al., 2024b; Chen et al., 2025). As a result, the LLM produces a diverse reasoning tree rather than a single chain.
PRMs are primarily categorized as either discriminative or generative (Malik et al., 2025). A discriminative PRM functions as a sequence classifier; in a single forward pass, it takes a full reasoning path as input and outputs a score for each intermediate step (He et al., 2024; Wang et al., 2023; Lightman et al., 2023). In contrast, a generative PRM is an auto-regressive model that must first generate its own textual critique before providing a final score, a significantly more expensive process (Zhao et al., 2025). Due to their superior balance of model parameters, reasoning quality, and hardware efficiency, discriminative PRMs are the preferred choice for state-of-the-art, verifier-guided TTS systems, particularly for memory-constrained edge deployment (Malik et al., 2025). Hence, our system focuses on discriminative PRMs. In contrast, as noted by (Snell et al., 2024; Liu et al., 2025), multi-step lookahead approaches (Chen and Li, 2024; Wang et al., 2025a), such as Monte Carlo Tree Search (MCTS) (Feng et al., 2023), introduce significant sampling and latency overhead with inferior accuracy, hence we do not consider them in this work.
@@ 2.3. LLM Serving
Serving frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024a) have been developed to optimize throughput and latency in streaming query scenarios. These systems incorporate key optimizations, including KV cache management to avoid recomputing attention states, paged attention to reduce GPU memory fragmentation, and preemptive scheduling to handle memory constraints by swapping requests. For TTS serving, goodput will be a more useful metric rather than throughput, as not all generated tokens will be selected for the final output. Despite its importance, no serving system to date natively supports the structured, multi-path search required for TTS in reasoning tasks.
@@ 3.1. Patterns Analysis in TTS Methods
Recent advancements in LLM reasoning have led to a variety of TTS methods, evolving from simple parallel sampling to more sophisticated search methods (Zhang et al., 2025; Chen et al., 2025; Snell et al., 2024; Hooper et al., 2025; Beeching et al., ). As illustrated in Fig. 2, this evolution marks a structural shift in the generation process: from exploring parallel but independent chains (e.g., CoT and Best-of-N) to constructing structured reasoning trees that allow for intermediate guidance and pruning (e.g., Beam Search, MCTS, and their variants). While recent studies show that small models can attain strong reasoning ability via long sequential CoT (Team and others, 2025), tree-structured reasoning remains essential in edge settings. First, it is orthogonal to sequential CoT and can further enhance the performance of small models. Second, parallel tree-based search has been demonstrated to be substantially more latency-efficient and token-efficient than purely sequential scaling (Wang et al., 2025b; Ghosal et al., 2025), which is critical for maintaining responsiveness under the tight resource constraints of edge deployment.
@@
This two-step process repeats until all reasoning paths reach a terminal state. Various search algorithms shown in Fig. 2 can be understood as specific implementations of this general pattern, differing in the heuristics they apply during the Generation or Verification stage. For instance, during Verification, standard Beam Search selects the top-K candidates globally with a static branching factor. In contrast, Diverse Selection (Snell et al., 2024; Beeching et al., ) modifies this to improve diversity by choosing the top candidate from distinct subtrees, while Dynamic Branching (Hooper et al., 2025; Wu et al., 2024b) makes the branching factor itself adaptive to verifier scores and system state. Other methods, like VG-Search (Chen et al., 2025), instead modify the Generation stage by altering the length of the thinking steps with varying verification granularities.
To understand the accuracy–latency trade-offs across different TTS methods, we conduct evaluations on the MATH-500 dataset. As illustrated in Fig. 3 (left), while advanced search methods often achieve higher algorithm accuracy, their overall latency remains a critical bottleneck. To address this system-level performance gap, we analyze the challenges shared by the abstracted TTS pattern.
@@ 3.2.3. Constrained Memory for Multi-Model Execution
4. FastTTS: Method and Optimization
@@ 4.1.2. Two-Phase Scheduling with Preemption
The speculative phase is fully preemptible: if a new request arrives or a running request is preempted due to memory constraints, all speculative generation is immediately stopped, and the system reverts to Phase 1 to serve the new request. This two-phase design ensures minimal overhead and preserves low-latency responsiveness. Crucially, the speculative phase does not introduce extra tail latency, as all speculative executions are strictly terminated—regardless of their progress—once all the standard beam generations for the current step finish.
@@ 4.3.1. Roofline-Guided KV Allocation
$T_{\mathrm{roof}}=\max\!\left(\frac{\mathrm{FLOPs}}{P},\;\frac{\mathrm{Bytes}}{\mathrm{BW}}\right),$
where $P$ is the device’s peak compute and $\mathrm{BW}$ its peak memory bandwidth, per hardware specification.
@@ 4.3.2. Extended Search Space with Offloading
This incurs a transfer overhead $T_{\text{overhead}}^{\text{offload}}$. The system then selects the lower-latency strategy: i) the optimal execution time $T_{\mathrm{tot}}$ from allocation search under the original constraint, or ii) the offloading time $T_{\mathrm{tot}}^{\text{offload}}+T_{\text{overhead}}^{\text{offload}}$, where $T_{\mathrm{tot}}^{\text{offload}}$ is computed from the maximum batch sizes allowed by the relaxed constraints. This dual-strategy policy lets FastTTS always pick the better option.
@@ 5. Implementation
FastTTS is implemented in ~6,500 lines of Python on top of vLLM (v0.9.2), operating the generator and verifier in separate worker processes via Python’s multiprocessing library. We extend the core LLMEngine of vLLM to implement our two-phase, preemptive scheduling policy, which dynamically switches between continuous batching and Speculative Beam Extension based on the request queue status. For the generator, we extend the default scheduler with our Dynamic Prefix-Aware Scheduling that implements a greedy heuristic to group beams from the same parent, maximizing KV cache reuse. Our Asymmetric Multi-Model Memory Allocation policy is managed by a lightweight searcher that is invoked dynamically to determine the partition of the KV cache between workers. The system exposes a configurable interface for various TTS strategies and hyperparameters.
Figure 11. Precise Goodput improvement of FastTTS over the vLLM baseline across different search algorithm variants. Experiments use the 1.5B+1.5B configuration on AIME. In dynamic branching, each beam branches proportionally to its verifier score; in varying granularity, the maximum step length is 64 tokens for the first 3 steps and 2048 thereafter.
Figure 12. FastTTS Goodput Improvement. The $x$-axis represents the number of beams (n).
@@ 6.1. Experimental Setup
Models. To assess FastTTS under diverse workloads, we evaluate three generator-verifier configurations designed to stress different system aspects, following common practice in prior work (Snell et al., 2024; Chen et al., 2025; Beeching et al., ; Liu et al., 2025). We test a verifier-heavy setting (1.5B+7B: Qwen2.5-Math-1.5B generator with a Math-Shepherd-Mistral-7B verifier) and a generator-heavy setting (7B+1.5B: Qwen2.5-Math-7B generator with a Skywork-o1-Open-PRM-1.5B verifier), both allocate 90% of GPU memory to test throughput limits. To simulate a highly resource-limited environment, we also test a memory-constrained setting (1.5B+1.5B: a 1.5B generator and verifier) (Yang et al., 2024; He et al., 2024; Wang et al., 2023), restricting it to 40% of GPU memory.
@@
Baseline Implementation. Our baseline system is built on top of the widely-used vLLM framework (version 0.9.2). We implement a standard verifier-guided test-time search, running the generator and verifier as separate vLLM instances, with remaining details following Hugging Face’s official search-and-learn implementation (Beeching et al., ). This baseline represents a naive but robust implementation of TTS, against which we compare the performance gains achieved by the optimizations in FastTTS.
@@
  • Completion Time: We measure the average end-to-end time taken per completion for a problem.